### v2

# Linear scaling causal discovery from high-dimensional time series by dynamical community detection

Abstract
Understanding which parts of a dynamical system cause each other is extremely relevant in fundamental and applied sciences. However, inferring causal links from observational data, namely without direct manipulations of the system, is still computationally challenging, especially if the data are high-dimensional. In this study we introduce a framework for constructing causal graphs from high-dimensional time series, whose computational cost scales linearly with the number of variables. The approach is based on the automatic identification of dynamical communities, groups of variables which mutually influence each other and can therefore be described as a single node in a causal graph. These communities are efficiently identified by optimizing the Information Imbalance, a statistical quantity that assigns a weight to each putative causal variable based on its information content relative to a target variable. The communities are then ordered starting from the fully autonomous ones, whose evolution is independent from all the others, to those that are progressively dependent on other communities, building in this manner a community causal graph. We demonstrate the computational efficiency and the accuracy of our approach on discrete-time and continuous-time dynamical systems including up to 80 variables.
preprint: APS/123-QED
Introduction — The growing abundance of time series - spanning fields from environmental monitoring to finance and neuroscience - offers unprecedented opportunities for understanding real world observations. Specifically, causal discovery allows studying how different parts of a system influence each other and inferring the existence of directional couplings (causal relationships) between variables in a time series [1, 2, 3].
Relationships between variables can be depicted, using Pearl’s approach [4, 5], as a time series graph. Each dynamic variable at a specific time is represented by a node and an arrow from one node to another represents a direct causal link. Such a graph encodes all conditional independence relationships between pairs of variables [6, 7, 2]. Several strategies can be used to build this graph. A first option involves checking for each couple of lagged variables if there is no conditioning set that makes them independent. An arrow between the two nodes is then drawn only if such set does not exist [8]. A strategy to assess this condition is to employ an iterative approach, where the size of the tested set is progressively increased [9, 10]. Although this approach leads to optimal detection power, it requires performing a number of tests that scales exponentially with the number of variables. Therefore, strategies to reduce the search space in practical applications have been developed [11, 12]. Alternatively, one can consider for each pair of lagged variables the largest possible conditioning set, which includes all past history of the time series. The multivariate versions of Granger Causality [13, 14, 15] and Transfer Entropy [16] can be regarded as implementations of this second strategy [8, 2]. Although this approach drastically reduces the number of tests to be performed, requiring a single test per pair of variables, it is significantly affected by the curse of dimensionality, often resulting in low detection power [11].
In this work, we introduce a method for causal discovery designed for high-dimensional time series. A powerful feature of our approach is that it enables the identification of causal influences that emerge from the collective dynamics of multiple variables at a very moderate computational cost, which, importantly, scales linearly with the number of variables. The graph obtained with our algorithm is a “mesoscopic” version of the standard graphs, as it groups together variables whose evolution cannot be described independently. We call these groups of variables dynamical communities. Our approach aims at revealing, if present, a hierarchical organization of these communities, emerging from unidirectional inter-group interactions. We refer to the resulting graph as community causal graph.
The key ingredient of our method is the efficient identification of the dynamical communities. This is achieved by optimizing the Information Imbalance [17, 18] as a function of a set of variational weights, one for each variable of the system, according to a prediction criterion broadly inspired by Granger Causality [14, 15]. We will show that the values of the weights allow identifying efficiently the communities and, as a consequence, building a causal graph, avoiding combinatorial searches of conditioning sets.
We build the method on the assumption of causal sufficiency, namely that there are no unobserved common drivers of two or more dynamic variables. In the Supplemental Material (SM) we discuss the effect of unobserved variables in different scenarios. We demonstrate the effectiveness of our approach using time series generated from both discrete-time and continuous dynamical systems, showcasing its applicability to high-dimensional scenarios.
The algorithm exploits the Information Imbalance [17], which allows quantifying the information content of different distance measures defined on a data set. The underlying idea is that a distance measure $d^{A}$ is predictive with respect to another distance measure $d^{B}$ if points close according to $d_{A}$ are also close according to $d^{B}$. The Information Imbalance is defined as
$\Delta(d^{A}\rightarrow d^{B})=\frac{2}{N^{2}}\sum_{\underset{(i\neq j)}{i,j}}\delta_{r^{A}_{ij},1}r^{B}_{ij}\,,$ (1)
where $N$ is the number of points in the dataset, $\delta$ is the Kronecker delta and $r_{ij}^{(\cdot)}$ is the distance rank of point $j$ with respect to point $i$. The superscript refers to the distance used in the computation. For example, $r_{ij}^{A}=2$ if $j$ is the second nearest neighbor of $i$ according to distance $d^{A}$. Eq. (1) defines a quantity that, in the limit of large $N$, approaches zero when all nearest neighbors in space $A$ remain nearest neighbors in space $B$, namely when $d^{A}$ is maximally predictive of $d^{B}$.
In ref. [19] we showed that the Information Imbalance can be used to infer the presence of causality between two multidimensional dynamical systems $X$ and $Y$. We assumed that if $X$ causes $Y$ and one attempts to make a prediction of the future of $Y$, a distance measure built using the present states of both $X$ and $Y$ will have more predictive power than a distance built using only $Y$. Formally, we assumed that $X$ causes $Y$ if and only if
$\hat{w}=\operatorname*{argmin}_{w}\Delta\big{(}d^{wX(0),Y(0)}\rightarrow d^{Y\left(\tau\right)}\big{)}\neq 0$ (2)
for some positive time lag $\tau$. The notation $d^{(\cdot)}$ denotes the squared Euclidean distance built over the superscript variables. For example,
$d_{\,ij}^{wX(0),Y(0)}=w^{2}\|X_{i}(0)-X_{j}(0)\|^{2}+\|Y_{i}(0)-Y_{j}(0)\|^{2}\,,$ (3)
where the Latin letters $i$ and $j$ denote independent realizations of the same dynamics, obtained either from uncorrelated samples of a single stationary trajectory, or from distinct trajectories with independent initial conditions. In the SM we relate the criterion in Eq. (2) to the notion of conditional independence. Importantly, the approach in ref. [19] builds on prior knowledge of the groups of variables that make up the distinct dynamical systems interacting with each other (the dynamical communities in the language of this work, see below). This is a very strong assumption, which for real-world data is typically violated. This work is dedicated to overcoming this problem.
Here, we extend the approach described above to automatically and efficiently find those dynamical communities. Our algorithm makes use of a differentiable extension of the Information Imbalance (DII) [18]:
$\text{DII}\,(d^{A}\rightarrow d^{B})=\frac{2}{N^{2}}\sum_{\underset{(i\neq j)}{i,j}}\frac{e^{-d^{A}_{ij}/\lambda}}{\sum_{m(\neq i)}e^{-d^{A}_{im}/\lambda}}r_{ij}^{B}\,.$ (4)
We note that Eq.(4) tends to Eq. (1) in the limit $\lambda~{}\rightarrow~{}0$. If distance $d^{A}$ depends on a set of parameters $\bm{w}$, this formulation allows optimizing such parameters by gradient descent.
Our algorithm, illustrated in Fig. 1, can be conceptually divided into three parts: in part i) the DII is minimized to infer each variable’s autonomous set, namely the set of variables that directly or indirectly cause it; in part ii), the autonomous sets are used to find the dynamical communities, namely sets of variables which directly or indirectly influence each other; and in part iii), a macroscopic graph depicting the causal interactions between the communities is constructed. The input to the algorithm is a set of $D$ time-dependent variables $\{X^{\alpha}(t)\}_{\alpha=1}^{D}$. We will refer to such variables as “microscopic”, in contrast to the “mesoscopic” dynamical communities identified by the method.
Figure 1: Illustration of the algorithm, using a 15-dimensional dynamical system composed of three groups of noisy coupled logistic maps. On the left, the ground-truth connectivity of the system is depicted in an all-variable representation. Repeated optimizations of DII$(d^{\bm{w}\odot\bm{X}(0)}\rightarrow d^{X^{\beta}(\tau)})$ are carried out for each target variable $\beta$ according to Eq. (5) ($\beta=0,5$ and 10 are shown in blue, red and green, respectively). The optimal weights from each optimization, depicted in the central bar plots, are used to construct a connectivity matrix $G$ (here, $G$ is built using a single time lag $\tau=1$). From this matrix, dynamical communities are identified and depicted on a graph, where several variables are grouped within the same node.
i) Identification of autonomous sets — As a first step we infer the autonomous set $\mathcal{S}^{\beta}$ of each variable $X^{\beta}$, which we define as the set of all variables $\{X^{\alpha}\}$ directly or indirectly causing $X^{\beta}$. $X^{\alpha}$ is a direct cause of $X^{\beta}$ if, for some time lag $\tau$, there exists a direct link $X^{\alpha}(0)\rightarrow X^{\beta}(\tau)$ in the ground-truth time series graph. Conversely, $X^{\alpha}$ is an indirect cause of $X^{\beta}$ if, for any $\tau>0$, the directed paths connecting $X^{\alpha}(0)$ and $X^{\beta}(\tau)$ pass through at least a third variable $X^{\gamma}(\tau^{\prime})$ ($\gamma\neq\alpha$, $\gamma\neq\beta$, $0<\tau^{\prime}<\tau$). We will use the notations $X^{\alpha}\rightarrow X^{\beta}$ and $X^{\alpha}\in\mathcal{S}^{\beta}$ interchangeably, without distinguishing between direct and indirect links among the microscopic variables.
To identify the autonomous sets $\mathcal{S}^{\beta}$, we optimize the Information Imbalance of Eq. (4) between a distance measure built with all dynamic variables at time $t=0$ and a distance built with a single variable $X^{\beta}$ at time $t=\tau$. To level out the fluctuation ranges of different variables, we first scale each variable by its standard deviation over the entire trajectory. Thus, the weights obtained by this optimization are
$\hat{\bm{w}}_{\beta}=\operatorname*{argmin}_{\bm{w}}\ \text{DII}\left(d^{\bm{w}\odot\bm{X}(0)}\rightarrow d^{X^{\beta}(\tau)}\right)\,,$ (5)
where $\odot$ denotes the element-wise product. We generalize the principle of Eq. (2) by stating that $X^{\alpha}$ is a direct or indirect cause of $X^{\beta}$ when the $\alpha$-component of $\hat{\bm{w}}_{\beta}$, denoted by $\hat{w}_{\beta}^{\alpha}$, is nonzero.
Following the intuition that different couplings might manifest at different time scales [20, 21, 19] we repeat the optimization in Eq. (5) for several values of $\tau$ between 1 and $\tau_{\text{max}}$, where $\tau_{\text{max}}$ is a hyper parameter which, in applications, we take of the order of the autocorrelation time of $X^{\beta}$. As depicted in Fig. 1, the maximum weights over the tested values of $\tau$ are stored as columns of a $D\times D$ matrix $G$:
$G^{\alpha\beta}=\max_{\tau}\hat{w}_{\beta}^{\alpha}\,.$ (6)
Constructing $d^{A}$ over a single time frame provides the correct results when the lag of direct links is not larger than 1 in the underlying time series graph. In the SM we describe how our approach can be extended when this condition does not hold, constructing $d^{A}$ on multiple time frames.
Ideally, each autonomous set $\mathcal{S}^{\beta}$ can be directly extracted from the nonzero elements of $G^{\alpha\beta}$. However, in applications $G$ is estimated using a finite number of measures (or a finite trajectory). Therefore, one can decide if an element is zero only according to a specified tolerance: we then set $X^{\alpha}\in\mathcal{S}^{\beta}$ whenever $G^{\alpha\beta}>\varepsilon$. The threshold $\varepsilon$ is the main hyper parameter of our algorithm. A false negative may appear if the coupling that we aim to detect by testing $\hat{w}_{\beta}^{\alpha}>0$ is present but weak. This may occur, for example, when $X^{\alpha}$ causes $X^{\beta}$ indirectly via several mediating variables.
To account for false negatives, we first construct a directed graph represented by the set of links for which $G^{\alpha\beta}>\varepsilon$, and then we construct each autonomous set $\mathcal{S}^{\beta}$ as the full set of ancestors of $X^{\beta}$. Missing links $X^{\alpha}\rightarrow X^{\beta}$ are likely to be recovered through indirect paths $X^{\alpha}\rightarrow X^{\gamma}\rightarrow...\rightarrow X^{\beta}$.
ii) Identification of dynamical communities — A hierarchical structure of dependencies between groups of variables can be directly retrieved from the analysis of the sets $\mathcal{S}^{\beta}$.
In particular, given an autonomous set $S^{\beta}$, we will define it as minimal if, for every variable $x^{\alpha}\in S^{\beta}$, one has $S^{\alpha}\equiv S^{\beta}$. In a minimal autonomous set, each variable depends on all and only the other variables in the set: together they form a dynamical community $\mathcal{G}_{k}$ whose evolution is independent of the rest of the network. Once all sets $\mathcal{G}_{k}$ have been identified, their variables can be removed from the graph, and new sets with the same property can be extracted. This will reveal new dynamical communities, which we distinguish from the initial ones by assigning them a higher autonomy level. By convention, we assign autonomy level 0 to the first dynamical communities identified, and we increase such a level by 1 at each iteration. Notice that, if a group has level of autonomy $k$, then it can be caused only by groups with level $m<k$, and, in turn, it can cause only groups with level $m>k$; no links can be present among groups at the same level.
Figure 2: Outcome and performance of the algorithm in three different test cases: (a)-(c) 15 coupled logistic maps, (d)-(f) 5 coupled Lorenz systems and (g)-(i) 2 Lorenz 96 systems. Left panels (a),(d) and (g): community causal graphs produced by the algorithm, which correctly retrieves the ground-truth connectivity. Central panels (b), (e) and (h): all-variable graphs obtained from condition $G^{\alpha\beta}>\varepsilon$, setting $\varepsilon$ to 0.38, 0.6 and 3.5, respectively. Nodes in community and all-variable graphs are colored according to the autonomous levels identified in step ii). Right panels (c), (f) and (i): validation measures of the recovered connectivity as a function of the threshold parameter $\varepsilon$, in the range $\left[0,\max_{\alpha\beta\,(\alpha\neq\beta)}G^{\alpha\beta}\right]$. For larger values of $\varepsilon$, no pairs of variables are found to be linked. Blue solid curve: adjusted mutual information between the retrieved groups and the ground-truth ones. Red dashed curve: accuracy of the all-variable adjacency matrix recovered from the final community graph, defined as the fraction of correctly identified links. Black star marker: results by setting the threshold $\varepsilon$ to the average of all weights in $G^{\alpha\beta}$.
iii) Construction of the community causal graph — The causal connections between the dynamical communities are depicted using a directed acyclic graph where each node represents a dynamical community $\mathcal{G}_{k}$, which we name community causal graph (see Fig. 1). We draw a directed arrow $\mathcal{G}_{k}\rightarrow\mathcal{G}_{m}$ if $l(\mathcal{G}_{m})=l(\mathcal{G}_{k})+1$ and if there is at least a pair of variables $X^{\alpha}\in\mathcal{G}_{k}$ and $X^{\beta}\in\mathcal{G}_{m}$ such that $X^{\alpha}\rightarrow X^{\beta}$.
By construction, links that connect communities with consecutive levels of autonomy are direct (e.g., $\mathcal{G}^{k}\rightarrow\mathcal{G}^{m}$), whereas links between non-consecutive communities may be indirect. In Appendix A, we show that one can distinguish between direct and indirect links among non-consecutive communities by a simple a posteriori analysis.
Results — We tested our approach on trajectories generated by dynamical systems of different complexity. Specifically, we considered three groups of five coupled logistic maps (Figs. 2a-c), five coupled Lorenz oscillators [22] (Figs. 2d-f), and two 40-dimensional Lorenz 96 systems [23] unidirectionally coupled (Figs. 2g-i). Since deterministic relationships are known to violate a condition known as faithfulness [2, 8] we added a small white noise to each variable while integrating the dynamic equations (see Appendix B). We report in the SM additional tests in presence of observational noise. For each system, distances entering the DII optimization at step i) of the algorithm were computed by extracting $N=2000$ frames, used as independent initial conditions, from a single time series realization.
The left panels in Fig. 2 show the community graphs produced by the algorithm when the threshold $\varepsilon$ is set to the average value of the weights in matrix $G$ (black star marker in the right panels). These graphs reproduce the correct connectivity read from the ground-truth equations: all variables grouped in a single node are dynamically intertwined, namely each one is a direct or indirect cause of any other variable, and an arrow between two groups at consecutive levels is present when at least two variables, one per group, are interacting. The central panels show the all-variable graphs built from the matrix $G^{\alpha\beta}>\varepsilon$, with nodes colored according to the level of autonomy assigned to the corresponding groups at step ii). This matrix was computed according to Eq. (6), with $\tau_{\text{max}}=1$ for the top row example, $\tau_{\text{max}}=60$ for the middle row, and $\tau_{\text{max}}=30$ for the bottom row example.
The outcome of the algorithm is influenced by the choice of the threshold $\varepsilon$. In Fig. 2, right column, we validated the correctness of the recovered dynamical communities and their connectivity separately, by measuring as a function of $\varepsilon$ the adjusted mutual information [24] between the true groups and the recollected ones (right panels, blue solid curves), and the accuracy of the all-variable adjacency matrix obtained from the final community graphs (right panels, dashed red curves). Both measures are defined in the range $[0,1]$, with 1 being the case of optimal reconstruction. We refer to the SM for a formal definition of these measures. Remarkably, in all systems we observe a wide range of threshold values leading to an exact reconstruction, and a false positive rate which is nonzero only for negligibly small values of the threshold ($\epsilon<10^{-16}$, $\epsilon<10^{-13}$ and $\epsilon<0.55$ for the systems in panels a, d and g, respectively).
Discussion — We have demonstrated that our approach can efficiently reconstruct the causal structure underlying high-dimensional dynamical systems, providing a coarse-grained visualization of the system’s causal connectivity, which has been a topic of growing interest [25]. The most relevant feature of our method is that the number of optimizations — analogous to conditional independence tests — scales linearly with the number of variables, and still the approach is able to capture “multibody” synergistic causal effects that are hard to detect by constructing conditioning sets of increasing size [2].
For comparison, we applied a state-of-the-art method for causal discovery on time series data, PCMCI [11], to the coupled Lorenz 96 systems. This algorithm efficiently reduces the search space for conditioning sets, at the cost of introducing some hyperparameters. In the SM, we show that the two 40-dimensional communities are correctly identified for certain hyperparameter combinations. However, the link among the communities appears hard to detect, as it results from few inter-community connections.
We notice one could consider analyzing the G matrix with graph clustering methods [26], rather than using it to identify communities. One could also employ this matrix to speed up the search of relevant groups of variables according to other criteria [27, 28, 29, 30], even in presence of bidirectional couplings. The approach can be improved by estimating the statistical confidence of each element in $G$, allowing the application of confidence-based thresholds, instead of a fixed $\varepsilon$. We consider the usage of multiple time delays $\tau$ in Eq. (6) as a valuable feature of our approach. In the SM (Fig. S2) we show that using only $\tau=1$ would significantly decrease the detection power of the algorithm.
The codes implementing our algorithm are available in the Python library DADApy [31].
Acknowledgments — This work was partially funded by NextGenerationEU through the Italian National Centre for HPC (Grant No. CN00000013). A.L. also acknowledges financial support by the region Friuli Venezia Giulia (project F53C22001770002 received by A.L.).